You are given the following information for Smashville, Inc.

Cost of goods sold: $ 184,000
Investment income: $ 1,600
Net sales: $ 387,000
Operating expense: $ 88,000
Interest expense: $ 7,400
Dividends: $ 6,000
Tax rate: 30 %
Current liabilities: $ 12,000
Cash: $ 21,000
Long-term debt: $ 32,000
Other assets: $ 40,000
Fixed assets: $ 125,000
Other liabilities: $ 5,000
Investments: $ 36,000
Operating assets: $ 64,000

During the year, Smashville, Inc., had 17,000 shares of stock outstanding and depreciation expense of $19,000. At the end of the year, Smashville stock sold for $42 per share. Calculate the price-book ratio, price-earnings ratio, and the price-cash flow ratio. (Do not round intermediate calculations. Round your answers to 2 decimal places.)
Answer #1

Price Book Ratio = Market price of share / Book value per share

Book value per share = Net worth of company / Shares outstanding

Net worth = (21000 + 40000 + 125000 + 36000 + 64000 - 12000 - 32000 - 5000) = 237000

Book value per share = 237000 / 17000 = 13.94

Price to Boom ratio = 42 / 13.94 = 3.01

Price earning ratio = Market price / Earning per share

Earnings before depreciation and taxes = (387000 + 1600 - 184000 - 88000 - 7400) = 109200

Earnings after tax = (109200 - 19000) * (1 - 0.3) = 63140

Earnings per share = 63140 / 17000 = 3.71

Price earnings ratio = 42 / 3.71 = 11.32

Price to cash flow ratio = Price / Cash flow per share

Cash flow = Earnings after tax + Depreciation = 63140 + 19000 = 82140

Cash flow per share = 82140 / 4.83

Price to cash flow = 42 / 4.83 =  8.70